OpenStudy (anonymous):

Subtract. (4x3 + 2x2 + 3x + 8) – (x3 – 5x2 + 11x – 9) 3x3 – 3x2 + 14x – 1 3x3 + 7x2 – 8x + 17 3x3 – 3x2 – 8x – 1 3x3 + 7x2 + 14x + 17

OpenStudy (zehanz):

Same as the other question: You can only add/subtract like terms: 4x³-x³=3x³ 2x²-5x²=... 3x-11x=... 8-(-9)=... It is not hard, just try it yourself!